i have 2 questions I have been confused on
1) A pharmacist needs 500 milliliters of 20% phenobarbital solution but has only 5% and 25% phenobarbital solutions available. find how many milliliters of each he should mix to get the desired solution.


2) Rabbits in a lab are to be kept on a strict daily diet that includes 30 grams of protein, 16 grams of fat and 24 grams of carbohydrates. the scientist has only three food mixes available with the following grams of nutrients per unit.

Protein fat carbohydrate
mix A 4, 6, 3
mix B 6, 1, 2
mix C 4, 1, 12

Find how many units of each mix are needed daily to meet each rabbits dietary need.

Answer by stanbon(75887)

Let amount of 5% solution be x;.05x of this is phenob.
Amount of 25% must be 500-x; 0.25 of this is phenob.
EQUATION:
Phenob + Phenob = 0.20(500)
0.05x+ 0.25(500-x)=100
5x+25(500-x)=10000
5x-25x=10000-12500
-20x=-2500
x= 125 (millileters of 5% solution)
500-x=375 (millileters of 25% solution)
Cheers,
Stan H.